Magnesium hydroxide is 54.87% oxygen by
mass. How much oxygen is in 260 g of the
compound?
Answer in units of g
How many moles of oxygen is this?
Answer in units of mol
1
Expert's answer
2012-11-09T02:38:46-0500
m(O) = 260 * 0.5487 = 142.662 g
n(O) = 142.662 g / 16 g/mol = 8.91 mol